Physics 115A xxxxx Fall 2013 Statistical Physics due 9:30 am Tuesday, Nov. 12

PROBLEM SET 5

Reading: Chapter 5 and 6 in Reif. Midterm: Tuesday November 5 in class. The test will cover all material through Chapter 5. The test is closed book. No notes, calculator or cell phones are allowed. This homework assignment is due a week after the midterm.

Even though this assignment is due after the midterm, the problems from Chapter 5 below would be useful to review before the midterm.

Problems:

Reif 5.22

Reif 5.24

Reif 5.26

Reif 6.1

Reif 6.8 Reif 6.10